Log Book Entry No.176

Days, Calendar Created: 15 Feb 2019 17:21:20 Days, Calendar Last Revised: 15 Feb 2019 21:21:05 Person JacobChambers Text PH4100 Photometry observation 2


  • Present JC, AD
  • Began cooling at 16:45 to -12 deg C (~90%)
  • Filter wheel set to Clear, Red, Green, Blue
  • Centred on Capella and took GPS-fix. Encountered problem: telescope believed it was -1h RA from Capella however this offset was constant.
  • Edit: This seemed to just be an issue with the coordinates read-out on the paddle.
  • Centred on M35 and took 1000x0.1s exposures to focus.
  • Then took 60s exposure- significant effects of moonlight on image- change target to M36 (RA/DEC: 05 37 13.8/34 09 04)
  • Light pollution significantly reduced suggesting moon was indeed the source.
  • took 60s and 90s exp in all filters
  • moved to (RA/DEC: 05 37 31/34 07 50) and repeated.
  • took dark and bias frames before finishing observation at 9:30pm.

Back to: LogBookAstronomy

Subject PH4100 Photometry observation 2
Status blank
Topic attachments
I Attachment History Action Size Date Who Comment
Compressed Zip archivezip 20190215.zip r1 manage 45707.8 K 15 Feb 2019 - 21:17 JacobChambers  
Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r3 - 15 Feb 2019 - JacobChambers

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback